### v1

@@ 3.2 Convergence guarantees for least squares and compressed sensing
Remark 3.4 (Replacing SR-LASSO with QBCP).
Corollary 3 also holds for weighted QCBP (see (13)). Indeed, [5, Theorem 7.12] is formulated for weighted SR-LASSO, an inspection of its proof reveals that it also holds for weighted QCBP, up to replacing $\|\boldsymbol{n}\|_{2}$ with the QCBP tuning parameter $\eta$ (see also [5, Theorem 7.4], which the proof of [5, Theorem 7.12] crucially relies on).
Remark 3.5 (From algebraic to exponential rates).
Analogous guarantees of algebraic convergence for uniform recovery, and with exponential decay rates, hold as well. To obtain them it is enough to combine Theorem 3.1 with [5, Theorem 7.13] and [5, Theorem 7.14] respectively. It is worth observing that the corresponding exponential rates are not dimension independent, though.
Remark 3.6 (From scalar- to vector-valued solution maps).
@@
(24) $\min\limits_{\boldsymbol{Z}\in\mathbb{C}^{N\times|V|}}\|\boldsymbol{\Psi}\boldsymbol{Z}-\boldsymbol{B}\|_{F}^{2}.$
@@
(25) $\min\limits_{\boldsymbol{Z}\in\mathbb{C}^{N\times|V|}}\lambda\|\boldsymbol{Z}\|_{1,1,\boldsymbol{w}}+\|\boldsymbol{\Psi}\boldsymbol{Z}-\boldsymbol{B}\|_{F},$
which falls under the category of vector-valued compressed sensing (and more generally, compressed sensing for Hilbert-valued functions); see also [5, Section 8.2.5].
@@ 4 Numerical experiments
In this section, we evaluate the efficacy of least squares and compressed sensing for approximating the (real-valued) parametric diffusion map $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\boldsymbol{y}))$ defined in (17). The primary objective of these experiments is to assess the accuracy of the method across various graph structures, sizes, and diffusion settings. As a performance metric, we use the Root Mean Square Error (RMSE), defined by
$\textnormal{RMSE}=\sqrt{\sum_{i=1}^{m_{\text{test}}}\frac{1}{m_{\text{test}}}(f_{v}(\boldsymbol{y}_{i}^{\text{test}})-\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}}))^{2}},$
where $f_{v}(\boldsymbol{y}_{i}^{\text{test}})$ is the actual value and $\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}})$ is the predicted value. In each of the following experiments, the test set will consist of 1000 pairs $(\boldsymbol{y}_{i}^{\text{test}},f(\boldsymbol{y}_{i}^{\text{test}}))$ where the $y_{i}$ are randomly sampled from the uniform distribution on $\mathcal{U}=[-1,1]^{d}$.
In Section 4.1 we first present our experiments on synthetic random graphs, more specifically, generated using the Stochastic Block Model (SBM), in order to produce graphs with a community structure [23]. Then, in Section 4.2 we replicate some of the same experiments on a real-world graph-structured dataset from Twitter. We implement the proposed method in Python using the NetworkX library for graph generation and manipulation [21], the NumPy library for matrix operations to solve the diffusion equation, and the equadratures [35] and SciPy [39] libraries for the polynomial approximation. The Python code is available at https://github.com/k-yoan/surrogate_graph_diffusion.
@@ 4.1 Synthetic graphs
For simplicity, we generate graphs from the SBM such that one graph always has the same number of nodes per community. In the subsequent sections, each experiment are repeated 10 times. The geometric mean of RMSE obtained from those rounds is shown in bold colored lines/curves, while the variance is shown with a shaded region around the mean, corresponding to the geometric standard deviation of the RMSE.
@@ 4.1.1 Proof of concept
First, we evaluate the accuracy of least squares and compressed sensing over $m$ sample points, for different dimensions and multi-index sets (total degree and hyperbolic cross, defined in (7) and (8), respectively). We choose the orders $n$ such that the cardinality of the multi-index sets is around the same for both bases ($n=8$ for the total degree set and $n=20$ for the hyperbolic cross). Figure 2 shows that when $d=3$, we attain a high level of accuracy. Another thing to note is that when the number of sample points $m$ is less than the cardinality of the basis (represented by the vertical grey dashed line on both plots), QCBP and weighted QCBP yield better results than least squares. This is expected, as least squares only works when the system is overdetermined: as we increase the number of sample points, the average RMSE given by the least squares method decreases and starts performing better than the other two methods.
@@
Figure 3 shows similar results with $d=6$. The way in which hyperparameters such as the multi-index set $\Lambda$, the dimension $d$ and the number of nodes $|V|$ affect the results is explored in the following subsections.
@@ 4.1.2 Impact of cardinality of basis on accuracy
In Figure 4, we fix the dimension $d=3$ and the number of sample points $m=350$, for the total degree set and the hyperbolic cross. For both bases, we observe the same pattern: as we increase the cardinality of the multi-index set (i.e., the number of coefficients in the polynomial expansion), the average RMSE lowers drastically. The same experiment is replicated in Figure 5 with $m=1000$, where we note the same correlation between the error and the size of the basis. The only major difference between the two experiments lies with the least-squares method: in Figure 4 (when $m=350$), the error increases sharply as the cardinality becomes large, whereas we only see a downward trend in Figure 5. This is once again expected, as the least-squares method does not perform well when the cardinality of the set is greater than the number of sample points $m$.
@@ 4.1.3 Impact of dimension on a fixed size graph
In the experiments, we test for both the total degree set and the hyperbolic cross. For $d=3$, $d=6$ and $d=10$, the cardinalities are 2925, 3003 and 3003, respectively, for the total degree. The results are shown in the left plots of Figures 6, 7 and 8. Similarly, for the hyperbolic cross, the cardinalities are 3143, 3119 and 3076 (for $d=3$, $d=6$ and $d=10$ respectively). The results are shown in the right plots of Figures 6, 7 and 8. The overall pattern we observe from the three figures is that a lower dimension leads to a better accuracy.
@@ 4.1.4 Impact of number of nodes per community for fixed dimension
We once again plot the average RMSE against the number of sample points, for graphs with 5, 10 and 15 nodes per community. We only show results for $d=3$, with the total degree set as a basis and the least-squares method, as we would see a similar pattern for other multi-index sets and other methods. In Figure 9, we can see that polynomial approximation yields better results when there are less nodes per community (i.e., when the graph is smaller). As we increase the number of sample points, the difference in performance seems to be more pronounced. The experiment was reproduced twice (with orders $n=6$ and $n=12$) to show that we can still achieve very good accuracy for larger graphs by simply increasing other hyperparameters, such as the order $n$.
@@
4.2 Validation on a real-world example: the Twitter dataset
To validate our theoretical results over a real-world example, we approximate $u_{2}$ (the solution of the diffusion equation at node 2) over the Twitter Interaction Network for the US Congress [14, 27]. This network represents the Twitter interaction network for the 117th United States Congress, both House of Representatives and Senate. The base data was collected via the Twitter’s API, then the empirical transmission probabilities (represented by weighted edges) were quantified according to the fraction of times one member retweeted, quote tweeted, replied to, or mentioned another member’s tweet. The number of nodes of the graph is $|V|=475$ and the number of edges is $|E|=13,289$. We detected 2 communities exploiting the fluid community detection algorithm [31], which contain, respectively, 110 and 365 nodes.
Figure 10: Left: Twitter dataset visualization. Communities are detected by means of the fluid community detection algorithm. Right: Average RMSE on the approximation of the diffusion process over the Twitter dataset as a function of the number of samples.
We computed the average RMSE on the approximation of $u_{2}$ over 10 runs, ranging the number of training sample points from 10 to 60 random sample points. The RMSE is computed over 100 test random sample points.
Experiments were run on Galileo G100 by CINECA, on a cluster node with CPU Intel CascadeLake 8260, with 24 cores, 2.4 GHz, 384GB RAM. The experiments substantially confirm the trend already observed in Section 4.1.1 of greater performance of least squares optimization, compared to the ones of QCBP, when the problem is overdetermined, and the ability of compressed sensing to recover good approximation in the underdetermined regime.
@@ 5 Conclusion
In this work, we have developed surrogate models based on sparse polynomial approximation for diffusion processes on graphs. To support the proposed methodology, we have shown convergence theorems for least squares and compressed sensing-based polynomial surrogates (see Corollaries 2 and 3, respectively), proving near-optimal algebraic rates as a function of the number of samples $m$. In Section 4, we validated our theoretical findings by investigating the impact of single hyperparameters within the Stochastic Block Model (discussed in Section 4.1) and showcasing the performance of least squares and compressed sensing-based sparse polynomial approximations on a real-life dataset, such as Twitter (see Section 4.2).
@@ Appendix A Supplementary information about multi-index sets
A multi-index set $\Lambda\subseteq\mathbb{N}_{0}^{d}$ is lower if the following holds for every $\boldsymbol{\nu}$,$\boldsymbol{\mu}\in\mathbb{N}_{0}^{d}$:
[…]
where the inequality is understood componentwise.
@@
Remark A.1.
The union of lower sets with cardinality $s$ is a hyperbolic cross $\Lambda_{s-1}^{\mathrm{HC}}$ of order $s-1$ (see [5, Proposition 2.5]). This emphasizes the connection between the hyperbolic cross index set and lower sets, and the importance of the hyperbolic cross in relation to sparse polynomial approximation, when the set $S$ is unknown.
@@ Appendix B Proof of Corollary 2
To prove Corollary 2 we need an auxiliary result. Specifically, we need an adaptation of [2, Theorem 6.1] to the finite-dimensional case.
Theorem B.1.
Let $10<\epsilon<$ 1, $\varrho$ be either the uniform or Chebyschev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}\sim_{\text{i.i.d}}\varrho$. Then, there exists a set $S\subseteq\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$ for any $f:\mathcal{U}\rightarrow\mathbb{C}$ that admits a holomorphic extension to a Bernstein polyellipse $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathbb{C}^{d}$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f}$ to the problem (9) is unique and satisfies
$\|f-\hat{f}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq C(\boldsymbol{\rho},p)\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty}.$
Proof B.2.
To prove the theorem it is enough to apply [2, Theorem 6.1] combined with [2, Remark 5.2]. We only need to prove that the support $S$ of the best $s$-term approximation to $f$ satisfies $S\subset\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$. Let =$\mathcal{F}=\{\boldsymbol{\nu}\in\mathbb{N}_{0}^{\mathbb{N}}:\|\boldsymbol{\nu}\|_{0}<\infty\}\subset\mathbb{N}_{0}^{\mathbb{N}}$.
@@
[…]
$=\int_{[-1,1]^{d}}f(x_{1},\ldots,x_{d})\prod_{j=1}^{d}\psi_{\nu_{j}}(x_{j})\int_{[-1,1]^{\mathbb{N}}}\prod_{j>d}\psi_{\nu_{j}}(x_{j})dx$
[…]
$=0$
For this reason, the best $s$-term approximation support $S$ can be chosen in $\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$.
The proof of Corollary 2 comes directly by combining Theorem B.1 with Theorem 3.1 considering the following setting:
  • $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$,
  • $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component)
  • consequently, $\mathcal{O}=\mathring{\mathcal{K}}$.
Acknowledgments
The authors would like to warmly thank Jason Bramburger and Iacopo Iacopini for the fruitful discussions. The authors acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support. Giuseppe Alessio D’Inverno acknowledges the support provided by the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS00000043 – CUP G93C22000610007) project. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.
Giuseppe Alessio D’Inverno is also member of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM).

### v2

@@ 3.2 Convergence guarantees for least squares and compressed sensing
Remark 1 (Replacing SR-LASSO with QBCP).
Corollary 3 also holds for weighted QCBP (see (13)). Indeed, [1, Theorem 7.12] is formulated for weighted SR-LASSO, an inspection of its proof reveals that it also holds for weighted QCBP, up to replacing $\|\boldsymbol{n}\|_{2}$ with the QCBP tuning parameter $\eta$ (see also [1, Theorem 7.4], which the proof of [1, Theorem 7.12] crucially relies on).
Remark 2 (From algebraic to exponential rates).
Analogous guarantees of algebraic convergence for uniform recovery, and with exponential decay rates, hold as well. To obtain them it is enough to combine Theorem 1 with [1, Theorem 7.13] and [1, Theorem 7.14] respectively. It is worth observing that the corresponding exponential rates are not dimension independent, though.
Remark 3 (From scalar- to vector-valued solution maps).
@@
$\min\limits_{\boldsymbol{Z}\in\mathbb{C}^{N\times|V|}}\|\boldsymbol{\Psi}\boldsymbol{Z}-\boldsymbol{B}\|_{F}^{2}.$ (25)
@@
$\min\limits_{\boldsymbol{Z}\in\mathbb{C}^{N\times|V|}}\lambda\|\boldsymbol{Z}\|_{1,1,\boldsymbol{w}}+\|\boldsymbol{\Psi}\boldsymbol{Z}-\boldsymbol{B}\|_{F},$ (26)
which falls under the category of vector-valued compressed sensing (and more generally, compressed sensing for Hilbert-valued functions); see also [1, Section 8.2.5].
@@ 4 Numerical experiments
In this section, we evaluate the efficacy of least squares and compressed sensing for approximating the (real-valued) parametric diffusion map $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\boldsymbol{y}))$ defined in (19). The primary objective of these experiments is to assess the accuracy of the method across various graph structures, sizes, and diffusion settings. As a performance metric, we use the Root Mean Square Error (RMSE), defined by $\textnormal{RMSE}=\sqrt{\sum_{i=1}^{m_{\text{test}}}\frac{1}{m_{\text{test}}}(f_{v}(\boldsymbol{y}_{i}^{\text{test}})-\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}}))^{2}},$ where $f_{v}(\boldsymbol{y}_{i}^{\text{test}})$ is the actual value and $\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}})$ is the predicted value. In each of the following experiments, the test set will consist of 1000 pairs $(\boldsymbol{y}_{i}^{\text{test}},f(\boldsymbol{y}_{i}^{\text{test}}))$ where the $y_{i}$ are randomly sampled from the uniform distribution on $\mathcal{U}=[-1,1]^{d}$. In this setup we adopt an edge-dependent diffusion, which is a special case of the time-and-edge-dependent diffusion case described by equation (18) in Example 1, where $h_{j}(t)\equiv 1$ for $j\in[d]$. Namely, the matrix $C$ in equation (2) is of the form
$\boldsymbol{C}(t,\boldsymbol{y})=\begin{pmatrix}\frac{y_{\sigma(1,1)}+1}{2}\cdot\mathds{1}&\cdots&\frac{y_{\sigma(1,K)}+1}{2}\cdot\mathds{1}\\ \vdots&\ddots&\vdots\\ \frac{y_{\sigma(K,1)}+1}{2}\cdot\mathds{1}&\cdots&\frac{y_{\sigma(K,K)}+1}{2}\cdot\mathds{1}\\ \end{pmatrix}$. (27)
In Section 4.1 we first present our experiments on synthetic random graphs, more specifically, generated using the SBM, in order to produce graphs with a community structure. Then, in Section 4.2.1 we replicate some of the same experiments on a real-world graph-structured dataset from Twitter. We implement the proposed method in Python using the NetworkX library for graph generation and manipulation [30], the NumPy library for matrix operations to solve the diffusion equation, and the equadratures [31] and SciPy libraries for the polynomial approximation. The Python code is available at https://github.com/k-yoan/surrogate_graph_diffusion.
@@ 4.1 Synthetic graphs
For simplicity, we generate graphs from the SBM such that one graph always has the same number of nodes per community, and the edge probabilities $p_{i,j}$ are always set to $1$ (when nodes $i$ and $j$ belong to the same community) or $0.04$ (when nodes $i$ and $j$ are in different communities). In the subsequent sections, each experiment are repeated 20 times. The geometric mean of RMSE obtained from those rounds is shown in bold colored lines/curves, while the variance is shown with a shaded region around the mean, corresponding to the geometric standard deviation of the RMSE.
@@ 4.1.1 Proof of concept
First, we evaluate the accuracy of least squares and compressed sensing over $m$ sample points, for different dimensions and multi-index sets (total degree and hyperbolic cross, defined in (7) and (8), respectively). We choose the orders $n$ such that the cardinality of the multi-index sets is around the same for both bases ($n=8$ for the total degree set and $n=20$ for the hyperbolic cross). Figure 2 shows that when $d=3$, we attain a high level of accuracy.
@@
Another thing to note is that when the number of sample points $m$ is less than the cardinality of the basis (represented by the vertical grey dashed line on both plots), QCBP and weighted QCBP yield better results than least squares. This is expected, as least squares only works when the system is overdetermined: as we increase the number of sample points, the average RMSE given by the least squares method decreases and starts performing better than the other two methods.
Figure 3 shows similar results with $d=6$.
@@
The way in which hyperparameters such as the multi-index set $\Lambda$, the dimension $d$ and the number of nodes $|V|$ affect the results is explored in the following subsections.
@@ 4.1.2 Impact of cardinality of basis on accuracy
In Figure 4, we fix the dimension $d=3$ and the number of sample points $m=350$, for the total degree set and the hyperbolic cross. For both bases, we observe the same pattern: as we increase the cardinality of the multi-index set (i.e., the number of coefficients in the polynomial expansion), the average RMSE lowers drastically. The same experiment is replicated in Figure 5 with $m=1000$, where we note the same correlation between the error and the size of the basis. The only major difference between the two experiments lies with the least-squares method: in Figure 4 (when $m=350$), the error increases sharply as the cardinality becomes large, whereas we only see a downward trend in Figure 5.
@@
This is once again expected, as the least-squares method does not perform well when the cardinality of the set is greater than the number of sample points $m$.
@@ 4.1.3 Impact of dimension on a fixed size graph
In the experiments, we test for both the total degree set and the hyperbolic cross. For $d=3$, $d=6$ and $d=10$, the cardinalities are 2925, 3003 and 3003, respectively, for the total degree. The results are shown in the left plots of Figures 6, 7 and 8.
@@
Similarly, for the hyperbolic cross, the cardinalities are 3143, 3119 and 3076 (for $d=3$, $d=6$ and $d=10$ respectively). The results are shown in the right plots of Figures 6, 7 and 8. The overall pattern we observe from the three figures is that a lower dimension leads to a better accuracy.
@@ 4.1.4 Impact of number of nodes per community for fixed dimension
We once again plot the average RMSE against the number of sample points, for graphs with 5, 10 and 15 nodes per community. We only show results for $d=3$, with the total degree set as a basis and the least-squares method, as we would see a similar pattern for other multi-index sets and other methods. In Figure 9, we can see that polynomial approximation yields better results when there are less nodes per community (i.e., when the graph is smaller).
@@
As we increase the number of sample points, the difference in performance seems to be more pronounced. The experiment was reproduced twice (with orders $n=6$ and $n=12$) to show that we can still achieve very good accuracy for larger graphs by simply increasing other hyperparameters, such as the order $n$.
4.2 Validation on real-world examples
To validate our theoretical results on real-world examples, we approximate $u_{2}$ (the solution of the diffusion equation at node 2) over datasets retrieved from Twitter and from Facebook, respectively. We find communities in each network using the fluid community detection algorithm [32].
4.2.1 The Twitter dataset
The Twitter network [33, 34] represents the Twitter interaction network for the 117th United States Congress, both House of Representatives and Senate. The base data was collected via the Twitter’s API, then the empirical transmission probabilities (represented by weighted edges) were quantified according to the fraction of times one member retweeted, quote tweeted, replied to, or mentioned another member’s tweet. The number of nodes of the graph is $|V|=475$ and the number of edges is $|E|=13,289$. We detected 2 communities exploiting the fluid community detection algorithm, which contain, respectively, 110 and 365 nodes.
Figure 10: Left: Average RMSE on the approximation of the diffusion process over the Twitter dataset as a function of the number of samples. Right: Average RMSE on the approximation of the diffusion process over the Facebook dataset as a function of the number of samples.
We computed the average RMSE on the approximation of $u_{2}$ over 5 runs, ranging the number of training sample points from 10 to 60 random sample points. The RMSE is computed over 1000 test random sample points.
The experiments substantially confirm the trend already observed in Section 4.1.1 of greater performance of least squares optimization, compared to the ones of QCBP, when the problem is overdetermined, and the ability of compressed sensing to recover good approximation in the underdetermined regime.
4.2.2 The Facebook dataset
This dataset consists of “circles” (or “friends lists”) from Facebook. Facebook data was collected from survey participants using a specific Facebook app [34, 35]. The size of this network is one order of magnitude larger than Twitter’s. The number of nodes of the graph is $|V|=4,039$ and the number of edges is $|E|=88,234$. We detected 2 communities employing the fluid community detection algorithm, which contain, respectively, 887 and 3152 nodes.
We computed the average RMSE on the approximation of $u_{2}$ over 5 runs, ranging the number of training sample points from 100 to 700 random sample points. The RMSE is computed over 1000 test random sample points. Although performances in terms of RMSE loss are worse, due to the greater size of the graph, the qualitative observations made in Section 4.2.1 remain valid for the Facebook network.
@@ 5 Conclusion
In this work, we have developed surrogate models based on sparse polynomial approximation for diffusion processes on graphs. To support the proposed methodology, we have shown convergence theorems for least squares and compressed sensing-based polynomial surrogates (see Corollaries 2 and 3, respectively), proving near-optimal algebraic rates as a function of the number of samples $m$. In Section 4, we validated our theoretical findings by investigating the impact of single hyperparameters within the Stochastic Block Model (discussed in Section 4.1) and showcasing the performance of least squares and compressed sensing-based sparse polynomial approximations on real-life datasets (Section 4.2).
@@
Acknowledgments
The authors warmly thank Jason Bramburger and Iacopo Iacopini for fruitful discussions. GAD acknowledges the support provided by the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS00000043 – CUP G93C22000610007) project. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them. GAD acknowledges the support of INdAM Gruppo Nazionale of Calcolo Scientifico (GNCS). SB acknowledges the support of NSERC through grant RGPIN-2020-06766 and FRQ - Nature et Technologies through grants 313276 and 359708.
@@ Appendix A Supplementary information about multi-index sets
A multi-index set $\Lambda\subseteq\mathbb{N}_{0}^{d}$ is lower if the following holds for every $\boldsymbol{\nu}$$\boldsymbol{\mu}\in\mathbb{N}_{0}^{d}$$(\boldsymbol{\nu}\in\Lambda\text{ and }\boldsymbol{\mu}\leq\boldsymbol{\nu})\Longrightarrow\boldsymbol{\mu}\in\Lambda,$ where the inequality is understood componentwise.
@@
Remark 4.
The union of lower sets with cardinality $s$ is a hyperbolic cross $\Lambda_{s-1}^{\mathrm{HC}}$ of order s1s-1 (see [1, Proposition 2.5]). This emphasizes the connection between the hyperbolic cross index set and lower sets, and the importance of the hyperbolic cross in relation to sparse polynomial approximation, when the set $S$ is unknown.
@@ Appendix B Proof of Corollary 2
To prove Corollary 2 we need an auxiliary result. Specifically, we need an adaptation of [26, Theorem 6.1] to the finite-dimensional case.
Theorem 2.
Let $0<\epsilon<1$, $\varrho$ be either the uniform or Chebyshev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}\sim_{\text{i.i.d}}\varrho$. Then, there exists a set $S\subseteq\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$ for each fixed $f:\mathcal{U}\rightarrow\mathbb{C}$ that admits a holomorphic extension to a Bernstein polyellipse $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathbb{C}^{d}$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f}$ to the problem (9) is unique and satisfies, for every $0<p<1$, the bound $\|f-\hat{f}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq C(\boldsymbol{\rho},p)\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty}$.
Proof.
To prove the theorem it is enough to apply [26, Theorem 6.1] combined with [26, Remark 5.2]. We only need to show that the support $S$ of the best $s$-term approximation to $f$ satisfies $S\subset\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$. Let $\mathcal{F}=\{\boldsymbol{\nu}\in\mathbb{N}_{0}^{\mathbb{N}}:\|\boldsymbol{\nu}\|_{0}<\infty\}\subset\mathbb{N}_{0}^{\mathbb{N}}$.
@@
$c_{\boldsymbol{\nu}}$ $=\int_{\mathcal{U}}f(\boldsymbol{x})\psi_{\boldsymbol{\nu}}(\boldsymbol{x})d\boldsymbol{x}=\int_{[-1,1]^{d}}f(\boldsymbol{x})\prod_{j=1}^{d}\psi_{\nu_{j}}(x_{j})\int_{[-1,1]^{\mathbb{N}}}\prod_{j>d}\psi_{\nu_{j}}(x_{j})dx$
$=\int_{[-1,1]^{d}}f(\boldsymbol{x})\prod_{j=1}^{d}\psi_{\nu_{j}}(x_{j})dx_{1}\cdots dx_{d}\cdot\prod_{j>d,j\neq k}\int_{-1}^{1}\psi_{\nu_{j}}(x_{j})dx_{j}\cdot\underbrace{\int_{-1}^{1}\psi_{\nu_{k}}(x_{k})dx_{k}}_{=0}=0$
Hence, the best $s$-term approximation support $S$ can be chosen in $\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$. ∎
The proof of Corollary 2 comes directly by combining Theorem 2 with Theorem 1 considering the following setting: $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$, $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component) and, consequently, $\mathcal{O}=\mathring{\mathcal{K}}$.